# Gapfree graphs and powers of edge ideals with linear quotients

### v1

@@ 5. Vertex duplication and linear quotients of powers
Consider the remaining case that $ab$ is incident to $x$. Without loss of generality, we may assume that $a=x$. In this case, it is clear that all edges incident to $a$ are larger than $cd$.
@@
We now come to the third main theorem in the present paper.
@@
For simplicity of the notation, for monomials $M$ and $N$, we shall write $M:N$ for $M/\operatorname{lcm}(M,N)$, which is the monomial generator of the colon ideal $(M):(N)$. Consider any two generators $A=u_{i}e_{j}>u_{k}e_{\ell}=B$ of $I^{q+1}$ in ${\mathcal{M}}^{(q+1)}$, and assume that they appear in the ordering ${\mathcal{M}}^{(q+1)}$ the first time in this representation; that is, $e_{j}$ and $e_{\ell}$ are the largest among all edges in the decomposition of $A$ and $B$, respectively. Set
@@
We can assume that $e_{i}\not=e_{\ell}$, i.e., $j\not=\ell$. Write $u_{i}=L_{1}\cdots L_{q}$ and $u_{k}=M_{1}\cdots M_{q}$, where $L_{i}$’s and $M_{j}$’s are edges in $G$, and set $e_{i}=ab,e_{\ell}=cd$. As observed before, $ab\succeq L_{t}$ and $cd\succeq M_{t}$ for all $t=1,\dots,q$. Also, $ab\succ cd$. We shall analyze the following two possibilities: $ab\cap cd=\emptyset$ and $ab\cap cd\not=\emptyset$.
@@
Suppose that there exists an edge $uv\in E$ such that $uv\mid Q$. Set
@@
Since $C$ contains $ab$, we have $C>u_{h}\cdot M_{q-1}=u_{k}\cdot ad=B$. Furthermore, $C:B=u_{t}:u_{h}$ is a variable which divides $uv\mid Q$. Hence, we can assume now that such an edge $uv$ does not exist. Particularly, it follows that $Q$ is the product of variables (with possibly repetitions) in an independent set of $G$.
@@
Case 2a. Consider the case where there exists $\theta\in\{\gamma+1,\dots,q\}$ such that $L_{\theta}\subseteq Z=\{z_{1},\dots,z_{\gamma},d,e\}$.
@@
Again, we have $C>B$. Also, $C:B=x_{1}:e$, so either $B=C$ is redundant or $C:B=x_{1}$ divides $Q$.
@@
Clearly, $u_{g}>B/M_{q-1}$ in ${\mathcal{M}}^{(q)}$. Furthermore, $u_{g}:B/M_{q-1}=x_{1}x_{2}:de$ (which is necessarily not 1). It follows that either $u_{g}=B/M_{q-1}$, which then implies that $B=u_{g}M_{q-1}$ is redundant, or there exists a generator $u_{h}>B/M_{q-1}$ of $I^{q}$ ($u_{h}$ could be $u_{g}$) such that $u_{h}:B/M_{q-1}$ is either $x_{1}$ or $x_{2}$, which then implies that $u_{h}M_{q-1}:B$ is either $x_{1}$ or $x_{2}$ that divides $Q$. This case is completed by showing that $u_{h}M_{q-1}>B$, which is done in the following Claim.
Claim. $u_{h}M_{q-1}>B$ in ${\mathcal{M}}^{(q+1)}$.
Proof of Claim. Since $u_{h}>B/M_{q-1}$ in ${\mathcal{M}}^{(q)}$, either $u_{h}$ contains an edge $f$ of $G$ such that $f\succ ad$ or $u_{h}=u_{h}^{\prime}\cdot ad$, where $u_{h}^{\prime}$ is a generator of $I^{q-1}$ and $u_{h}^{\prime}>B/(M_{q-1}\cdot ad)=M_{1}\cdot M_{q-2}\cdot M_{q}$ in ${\mathcal{M}}^{(q-1)}$. If $u_{h}$ contains $f\succ ad$, then $u_{h}M_{q-1}>B$ in ${\mathcal{M}}^{(q+1)}$ as claimed.
@@
Case 2b. Consider the case where there does not exist such $L_{\theta}\subseteq Z$. Since $L_{\gamma+1}\cdots L_{q}$ is contained in $M_{\gamma+1}\cdots M_{q-1}\cdot d\cdot e\cdot z_{1}\cdots z_{\gamma}$, and $\deg(L_{\gamma+1}\cdots L_{q})=\deg(M_{\gamma+1}\cdots M_{q-1})+2$, there must exist $\theta,\delta\in\{\gamma+1,\dots,q\}$ and $\sigma\in\{\gamma+1,\dots,q-1\}$ such that: $L_{\theta}\cap Z\not=\emptyset,L_{\delta}\cap Z\not=\emptyset$, and $M_{\sigma}$ covers (consists of) two vertices, one from each of the edges $L_{\theta}$ and $L_{\delta}$. For simplicity of notations, we shall assume that $\sigma=3$.
@@
Then, $C>B$, and $C:B=x_{1}:d$. Thus, either $B=C$ is redundant, or $C:B=x_{1}$, which divides $Q$.
@@
For $Q=x_{1}x_{2}$, we must also have that $y_{2}$ is $d$ or $e$. That means, either $x_{2}d$ or $x_{2}e$ is an edge in $G$. The same argument for this situation in Theorem 6.4 now applies to complete the proof.
@@
Since $Q=x_{1}x_{2}x_{3}$, this implies that $y_{3}$ is either $d$ or $e$. That is, either $x_{3}d$ or $x_{3}e$ is an edge in $G$. The same argument of Theorem 6.4 in this case also applies to complete the proof.
@@
Our proof does not currently work for the implications $I^{4}\rightarrow I^{5}$, $I^{5}\rightarrow I^{6}$ and $I^{6}\rightarrow I^{7}$.
@@
The proof of Theorem 6.4 applies for relatively manageable graphs to also work for the implications $I^{4}\rightarrow I^{5},I^{5}\rightarrow I^{6}$ and $I^{6}\rightarrow I^{7}$; for instance, anticycles, i.e., the complementary graphs of cycles. Particularly, by exhibiting a specific linear quotients order for $I(G)^{2}$, when $G$ is an anticycle, it follows that $I(G)^{q}$ has linear quotients for all $q\geq 2$. This is a main result of a recent preprint [4].

### v2

@@ 5. Pentagon together with one vertex
Consider the remaining case that $ab$ is incident to $x$. Without loss of generality, we may assume that $a=x$. In this case, it is clear that all edges incident to $a$ are larger than $cd$. ∎
@@
Note that the orders ${\mathcal{M}}^{(k)}$ given in Definitions 6.3 look similar to the orders $\mathcal{N}^{(k)}$ described in Section 4, but are different in nature. We now come to the last main result of the paper.
@@
Consider any two generators $A=u_{i}e_{j}>u_{k}e_{\ell}=B$ of $I^{q+1}$ in ${\mathcal{M}}^{(q+1)}$, and assume that they appear in the ordering ${\mathcal{M}}^{(q+1)}$ the first time in this representation; that is, $e_{j}$ and $e_{\ell}$ are the largest among all edges in the decomposition of $A$ and $B$, respectively. Set
@@
We can assume that $e_{j}\not=e_{\ell}$, i.e., $j\not=\ell$. Write $u_{i}=L_{1}\cdots L_{q}$ and $u_{k}=M_{1}\cdots M_{q}$, where $L_{i}$’s and $M_{j}$’s are edges in $G$, and set $e_{j}=ab,e_{\ell}=cd$. As observed before, $ab\succeq L_{t}$ and $cd\succeq M_{t}$ for all $t=1,\dots,q$. Also, $ab\succ cd$. We shall analyze the following two possibilities: $ab\cap cd=\emptyset$ and $ab\cap cd\not=\emptyset$.
@@
Suppose that there exists an edge $uv\in E$ such that $uv\mid Q$. Without loss of generality, assume that $M_{q-1}$ is the largest among $M_{1},\dots,M_{q-1}$. Set
@@
Then, $C:B=u_{t}:u_{h}$ is a variable which divides $uv\mid Q$. Hence, we are done in this case by the following Claim.
Claim 1. $C>u_{h}\cdot M_{q-1}=B$ in ${\mathcal{M}}^{(q+1)}$.
Proof of Claim 1. Since $u_{t}>u_{h}$, either $u_{t}$ contains an edge $f$ of $G$ for which $f\succ ad$ or $u_{t}=u_{t}^{\prime}\cdot ad$, where $u_{t}^{\prime}$ is a generator of $I^{q-1}$ and $u_{t}^{\prime}>u_{h}/ad=M_{1}\cdots M_{q-2}\cdot be.$ In the first case, due to the existence of $f$, we have $u_{t}\cdot M_{q-1}>u_{h}\cdot M_{q-1}=B$. Consider the case when $u_{t}=u_{t}^{\prime}\cdot ad$, where $u_{t}^{\prime}>(M_{1}\cdots M_{q-2})\cdot be=N$ in ${\mathcal{M}}^{(q-1)}$.
If $M_{q-1}>be$, then this implies that $u_{t}^{\prime}\cdot M_{q-1}>N\cdot M_{q-1}$ in ${\mathcal{M}}^{(q)}$, and so $C=u_{t}^{\prime}\cdot M_{q-1}\cdot ad>N\cdot M_{q-1}\cdot ad=u_{h}\cdot M_{q-1}=B$ in ${\mathcal{M}}^{(q+1)}$. If $M_{q-1}<be$, then either $u_{t}^{\prime}$ contains an edge $f^{\prime}$ of $G$ such that $f^{\prime}\succ be\succ M_{q-1}$, or $u_{t}^{\prime}=be\cdot u_{t}^{\prime\prime}$, where $u_{t}^{\prime\prime}>M_{1}\cdots M_{q-2}$ in ${\mathcal{M}}^{(q-2)}$. The existence of $f^{\prime}$ gives $u_{t}^{\prime}\cdot M_{q-1}>N\cdot M_{q-1}$ in ${\mathcal{M}}^{(q)}$ and, again, implies that $C=u_{t}^{\prime}\cdot M_{q-1}\cdot ad>B$. On the other hand, if $u_{t}^{\prime}=be\cdot u_{t}^{\prime\prime}$ as said, then since $M_{q-1}\succ M_{i}$ for all $i=1,\dots,q-2$, we have $u_{t}^{\prime\prime}\cdot M_{q-1}>(M_{1}\cdots M_{q-2})\cdot M_{q-1}$, which implies that $u_{t}^{\prime}\cdot M_{q-1}=u_{t}^{\prime\prime}\cdot M_{q-1}\cdot be>N\cdot M_{q-1}$. As before, it follows that $C>B$. The claim is proved.
We proceed by assuming now that an edge $uv$ such that $uv\mid Q$ does not exist. Particularly, it follows that $Q$ is the product of variables (with possibly repetitions) in an independent set of $G$.
@@
Case 2.1. Consider the case where there exists $\theta\in\{\gamma+1,\dots,q\}$ such that $L_{\theta}\subseteq Z=\{z_{1},\dots,z_{\gamma},d,e\}$.
@@
Again, we have $C>B$. Also, $C:B=x_{1}:e$, so either $B=C$, which contradicts $cd$ being the largest edge in the decomposition of $B$, or $C:B=x_{1}$ divides $Q$.
@@
Clearly, $u_{g}>B/M_{q-1}$ in ${\mathcal{M}}^{(q)}$. Furthermore, $u_{g}:B/M_{q-1}=x_{1}x_{2}:de$ (which is necessarily not 1). It follows that either $u_{g}=B/M_{q-1}$, which then implies that $B=u_{g}M_{q-1}$ which contradicts $cd$ being the largest edge in the decomposition of $B$, or there exists a generator $u_{h}>B/M_{q-1}$ of $I^{q}$ ($u_{h}$ could be $u_{g}$) such that $u_{h}:B/M_{q-1}$ is either $x_{1}$ or $x_{2}$, which then implies that $u_{h}M_{q-1}:B$ is either $x_{1}$ or $x_{2}$ that divides $Q$. This case is completed by showing that $u_{h}M_{q-1}>B$, which is done in the following Claim.
Claim 2. $u_{h}M_{q-1}>B$ in ${\mathcal{M}}^{(q+1)}$.
Proof of Claim 2. In Claim 1, we have seen a similar statement for a simpler situation. The proof of this claim goes in the same line of arguments (with more details). Since $u_{h}>B/M_{q-1}$ in ${\mathcal{M}}^{(q)}$, either $u_{h}$ contains an edge $f$ of $G$ such that $f\succ ad$ or $u_{h}=u_{h}^{\prime}\cdot ad$, where $u_{h}^{\prime}$ is a generator of $I^{q-1}$ and $u_{h}^{\prime}>B/(M_{q-1}\cdot ad)=M_{1}\cdots M_{q-2}\cdot M_{q}$ in ${\mathcal{M}}^{(q-1)}$. If $u_{h}$ contains $f\succ ad$, then $u_{h}M_{q-1}>B$ in ${\mathcal{M}}^{(q+1)}$ as claimed.
@@
Case 2.2. Consider the case where there does not exist such $L_{\theta}\subseteq Z$. Since $L_{\gamma+1}\cdots L_{q}$ is contained in $M_{\gamma+1}\cdots M_{q-1}\cdot d\cdot e\cdot z_{1}\cdots z_{\gamma}$, and $\deg(L_{\gamma+1}\cdots L_{q})=\deg(M_{\gamma+1}\cdots M_{q-1})+2$, there must exist $\theta,\delta\in\{\gamma+1,\dots,q\}$ and $\sigma\in\{\gamma+1,\dots,q-1\}$ such that: $L_{\theta}\cap Z\not=\emptyset,L_{\delta}\cap Z\not=\emptyset$, and $M_{\sigma}$ covers (consists of) two vertices, one from each of the edges $L_{\theta}$ and $L_{\delta}$. For simplicity of notations, we shall assume that $\sigma=3$.
@@
Then, $C>B$, and $C:B=x_{1}:d$. Thus, either $B=C$ is redundant, or $C:B=x_{1}$, which divides $Q$. ∎
@@
For $Q=x_{1}x_{2}$, we must also have that $y_{2}$ is $d$ or $e$. That means, either $x_{2}d$ or $x_{2}e$ is an edge in $G$. The same argument for this situation in Theorem 6.4 now applies to complete the proof. ∎
@@
Since $Q=x_{1}x_{2}x_{3}$, this implies that $y_{3}$ is either $d$ or $e$. That is, either $x_{3}d$ or $x_{3}e$ is an edge in $G$. The same argument of Theorem 6.4 in this case also applies to complete the proof. ∎
@@
Our proof does not currently work to show that $I^{k}$ has linear quotients implies $I^{k+1}$ has linear quotients for $k=4,5,6$ in general.
@@
For relatively manageable graphs, the techniques from the proof of Theorem 6.4 can be used to show that $I^{k}$ has linear quotients implies $I^{k+1}$ has linear quotients for $k=4,5,6$; for instance, if $G$ is an anticycle, i.e., the complementary graph of a cycle. In particular, by exhibiting a specific linear quotients order for $I(G)^{2}$, when $G$ is an anticycle, it follows that $I(G)^{q}$ has linear quotients for all $q\geq 2$. This is a main result of a recent preprint [5].